### v1

@@ 3 The connections between problem (1) and the min-max reformulation (9)
(12) $F(x):=\begin{cases}\frac{f(x)}{g(x)}+h(x),&\text{if $x\in\Omega\cap\mathcal{C}$},\\ +\infty,&\text{else.}\end{cases}$
Obviously, problem (1) can be rewritten as $\min\{F(x):x\in\mathbb{R}^{n}\}$. Moreover, we use $\widetilde{F}$ to denote the extended objective of the min-max problem (9), i.e.,
(13) $\widetilde{F}(x,c):=2cf(x)-c^{2}f(x)g(x)+h(x)+\iota_{\Omega\cap\mathcal{C}}(x).$
By a direct computation, we obtain that for all $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$,
(14) $F(x)-\widetilde{F}(x,c)=f(x)g(x)\left(\frac{1}{g(x)}-c\right)^{2}.$
@@
(16) $x^{\star}\in\arg\min\left\{\max_{c\in\mathbb{R}}\widetilde{F}(x,c):x\in\Omega\cap\mathcal{C}\right\}.$
Besides, in view of (15), we derive that $1/g(x^{\star})\in\arg\max\{\widetilde{F}(x^{\star},c):c\in\mathbb{R}\}$, which along with (16) indicates that $(x^{\star},1/g(x^{\star}))$ is a global min-max point of (1.7), according to Definition 2.3.
Conversely, if $(x^{\star},c_{\star})$ is a global min-max point of (9), then we derive from Definition 2.3 that $x^{\star}$ satisfies (16) . Thus, we conclude that $x^{\star}\in\arg\min\{F(x):x\in\Omega\cap\mathcal{C}\}$ from (16) and (15). This completes the proof.
We next establish the connection between the local minimizes of problem (1) and the local min-max points of problem (9).
Proposition 3.3.
@@
Proof 3.4.
We first prove that the statement that $(x^{\star},1/g(x^{\star}))$ is a local min-max point of problem (9) follows from the $x^{\star}$ being a local minimizer of problem (1). Since (15) indicates that $1/g(x^{\star})$ is a global maximizer of $\widetilde{F}(x^{\star},\cdot)$, it suffices to show that, for all $\epsilon>0$, the $x^{\star}$ is always a local minimizer of
(17) $\varphi_{\epsilon}(x):=\max\left\{\widetilde{F}(x,c):c\in\mathbb{R},~{}|c-1/g(x^{\star})|\leq\epsilon\right\},$
relative to $\Omega\cap\mathcal{C}$ according to Definition 2.4. Let $\epsilon>0$. Firstly, due to the continuity of $1/g$ around $x^{\star}\in\Omega$, there exists $\delta_{\epsilon}>0$ such that $|1/g(x)-1/g(x^{\star})|\leq\epsilon$ for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})$. In view of this, we deduce from the definition (17) that
(18) $\widetilde{F}(x,1/g(x))\leq\varphi_{\varepsilon}(x)\text{,~{}~{} for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})\text{. }$
@@
(19) $F(x^{\star})\leq F(x)\text{,~{}~{} for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon}^{\prime})\cap(\Omega\cap\mathcal{C})\text{. }$
@@
(20) $\varphi_{\epsilon}(x^{\star})\stackrel{{\scriptstyle{(I)}}}{{\leq}}\max_{c\in\mathbb{R}}\widetilde{F}(x^{\star},c)\stackrel{{\scriptstyle{(II)}}}{{=}}F(x^{\star})\stackrel{{\scriptstyle{(III)}}}{{\leq}}F(x)\stackrel{{\scriptstyle{(IV)}}}{{=}}\widetilde{F}(x,1/g(x))\stackrel{{\scriptstyle{(V)}}}{{\leq}}\varphi_{\varepsilon}(x),$
@@
Conversely, we let $(x^{\star},c_{\star})$ be a local min-max point of (9). Then, according to Definition 2.4, we know that $c_{\star}$ is a local maximizer of $\widetilde{F}(x^{\star},\cdot)$. Hence, we have $0=\nabla(\widetilde{F}(x^{\star},\cdot))(c_{\star})$, which leads to $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. In addition, also according to Definition 2.4, we know that $x^{\star}$ is a local minimizer of $\varphi_{\epsilon}$ with some $\epsilon>0$ relative to $\Omega\cap\mathcal{C}$. Thus, there exists some $\delta_{\epsilon}>0$ such that
(21) $\varphi_{\epsilon}(x^{\star})\leq\varphi_{\epsilon}(x),\text{\quad for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$.}$
@@
(22) $F(x^{\star})=\widetilde{F}(x^{\star},c_{\star})\leq\varphi_{\epsilon}(x^{\star}),$
where the first equation follows from (15) and the fact that $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. As to the right side of (21), it is clear that
(23) $\varphi_{\epsilon}(x)\leq\max\left\{\widetilde{F}(x,c):c\in\mathbb{R}\right\}=F(x),\quad\text{for all $x\in\Omega\cap\mathcal{C}$.}$
Together with (22) and (23), we deduce from (21) that $F(x^{\star})\leq F(x)$ for any $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$, which yields that $x^{\star}$ is a local minimizer of (1).
Now we study the relationship between stationary points of problems (1) and (9). To this end, for $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$, we let ^$\widehat{\partial}_{x}\widetilde{F}(x,c)$ and $\nabla_{c}\widetilde{F}(x,c)$ denote the partial Fréchet subdifferential of $\widetilde{F}$ with respect to $x$ and the partial derivative of $\widetilde{F}$ with respect to $c$, respectively.
Proposition 3.5.
@@
[…]
@@
Proof 3.6.
Let $(x^{\star},c_{\star})\in(\Omega\cap\mathcal{C})\times\mathbb{R}$ Due to the continuity of $g$ around $x^{\star}\in\Omega$, there exists $\delta>0$ such that $g(x)>0$ holds for all $x\in\mathcal{B}(x^{\star},\delta)$. Then, we have
(25) ^$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})=\widehat{\partial}\left(f/g+h+\iota_{\mathcal{C}}-\varphi\right)(x^{\star})=\widehat{\partial}(F-\varphi)(x^{\star}),$
@@
(26) $\varphi(x)=f(x)g(x)\left(\frac{1}{g(x)}-c_{\star}\right)^{2},\text{~{}~{}~{} for all $x\in\mathcal{B}(x^{\star},\delta)$}.$
@@
Conversely, we suppose that (24) holds. Then $0=\nabla_{c}\widetilde{F}(x^{\star},c_{\star})$ implies that $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. Plugging this into (26) and noting that $f$ and $g$ are non-negative, we deduce that $\varphi(x^{\star})=0$ is the minimum of $\varphi$ on $x\in\mathcal{B}(x^{\star},\delta)$. Hence, by invoking the generalized Fermat’s Rule, we derive $0\in\widehat{\partial}\varphi(x^{\star})$. With the help of [30, Theorem 3.1], we have
(27) ^$\widehat{\partial}(F-\varphi)(x^{\star})\subseteq\bigcap_{v^{\star}\in\widehat{\partial}\varphi(x^{\star})}(\widehat{\partial}F(x^{\star})-v^{\star})\subseteq\widehat{\partial}F(x^{\star})-\{0\}.$
Together with the first relation of (24), the equation (25), and the relation (27), we finally conclude that $0\in\widehat{\partial}F(x^{\star})$. This completes the proof.
@@
(28) $0\in\widehat{\partial}(c_{\star}f+\iota_{\mathcal{C}})\left(x^{\star}\right)-c_{\star}^{2}f(x^{\star})\partial g\left(x^{\star}\right)+\nabla h_{1}\left(x^{\star}\right)-\partial h_{2}\left(x^{\star}\right).$
@@
(29) ^$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})\overset{{(I)}}{=}\widehat{\partial}(2c_{\star}f-c_{\star}^{2}fg+h+\iota_{\mathcal{C}})(x^{\star})$
=$\overset{{(II)}}{=}\widehat{\partial}(c_{\star}f-c_{\star}^{2}f(x^{\star})g-h_{2}+\iota_{\mathcal{C}})(x^{\star})+\nabla\varphi_{2}(x^{\star})+\nabla h_{1}(x^{\star})$
@@
where ${(I)}$ follows from the definition of $\widetilde{F}$ (see (13)) and the open $\Omega$, ${(II)}$ holds because $\varphi_{2}$ and $h_{1}$ are differentiable at $x^{\star}$, and ${(III)}$ holds thanks to [30, Theorem 3.1], $c_{\star}^{2}f(x^{\star})\geq 0$, and the real-valued convexity of the functions $g$ and $h_{2}$. Especially, ${(III)}$ in (29) becomes an equality when $g$ and $h_{2}$ are differentiable at $x^{\star}$. Invoking (29) and Proposition 3.5, we then obtain the desired results.
Inspired by Corollary 3.7, we introduce the definition of a critical point of $F$.
Definition 3.9.
@@
To close this section, we make some remarks regarding Definition 3.9. In view of Corollary 3.7 and the generalized Fermat’s Rule, a local minimizer of $F$ must be a critical point of $F$. In particular, when both $g$ and $h_{2}$ are differentiable at $x^{\star}$, according to Corollary 3.7 we see that the statement that $x^{\star}$ is a critical point of $F$ coincides with $0\in\widehat{\partial}F(x^{\star})$.
@@ 4 An alternating maximization proximal descent algorithm
In this section, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max reformulation (9) and establish its subsequential convergence to a critical point of problem (1) under a mild assumption.
@@ 4.1 AMPDA and its interpretations
In this subsection, we present the AMPDA and then make some interpretations on this algorithm. To this end, we first introduce the notion of proximal operators. For a proper closed function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the proximal operator of $\varphi$ at $x\in\mathbb{R}^{n}$ is denoted by $\mathrm{prox}_{\varphi}(x)$ and defined as
(30) $\mathrm{prox}_{\varphi}(x):=\arg\min\left\{\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathbb{R}^{n}\right\}.$
The proximal operator $\mathrm{prox}_{\varphi}$ takes an $x\in\mathbb{R}^{n}$ and maps it into a subset of $\mathbb{R}^{n}$, which may be empty, a singleton or a subset with multiple vectors. If for a given $x\in\mathbb{R}^{n}$, the function $z\to\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}$ is coercive, then $\mathrm{prox}_{\varphi}(x)$ is nonempty (see, for example, [4, Theorem 6.4]). Moreover, if $\varphi$ is convex, then $\mathrm{prox}_{\varphi}(x)$ is a singleton for all $x\in\mathbb{R}^{n}$. In particular, let $f_{\mathcal{C}}:=f+\iota_{\mathcal{C}}$, then the proximal operator $\mathrm{prox}_{f_{\mathcal{C}}}(x)$ is nonempty for all $x\in\mathbb{R}^{n}$ due to $f_{\mathcal{C}}(x)\geq 0$. Let $Q:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\to(-\infty,+\infty]$ be defined by
(31) $Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2cf(x)+c^{2}f(x)(g^{*}(y)-\left\langle{x},{y}\right\rangle)+h_{1}(x)+h_{2}^{*}(z)-\left\langle{x},{z}\right\rangle$
@@
(32) $F(x)\overset{{(I)}}{=}\widetilde{F}(x,c_{x})\overset{{(II)}}{\leq}Q(x,y,z,c_{x}),$
where ${(I)}$ follows from (15), ${(II)}$ follows from the Fenchel-Young Inequality and the non-negativity of $c_{x}^{2}f(x)$, and ${(II)}$ becomes an equality when $(y,z)\in\partial g(x)\times\partial h_{2}(x)$. Now we present our AMPDA stated in Algorithm 1. Before proceeding, we provide some interpretations and observations on the AMPDA. This algorithm can be interpreted as an alternating iteration method for solving the min-max reformulation (9) as illustrated below.
@@
Step 0. Input $x^{0}\in\mathrm{dom}(F)$, $0<\underline{\alpha}\leq\overline{\alpha}$, $\sigma>0$, $10<\gamma<$ 1, and set $t\leftarrow 0$.
@@
(33) $\widehat{x}^{k}\in\mathrm{prox}_{\alpha c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right).$
@@
(34) $Q(\widehat{x}^{k},y^{k},z^{k},1/g(\widehat{x}^{k}))+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k});$
@@
Second, in Step 2 of the $k$-th iteration, AMPDA solves a proximal subproblem with respect to $f+\iota_{\mathcal{C}}$, which also requires to evaluate the gradient of $h_{1}$, the subgradients of $g$ and $h_{2}$ at $x^{k}$. As we will see below, this can be viewed as applying one step of the majorization-minimization scheme to the minimization problem $\min\{\widetilde{F}(x,c_{k}):x\in\mathbb{R}^{n}\}$. Since $f$ and $\nabla h_{1}$ are locally Lipschitz continuous on $\Omega$, there exists $\delta_{k}>0$ such that $f$ and $\nabla h_{1}$ are Lipschitz continuous on the neighborhood $\mathcal{B}(x^{k},\delta_{k})\subseteq\Omega$. Let $L_{f,k}>0$ and $L_{\nabla h_{1},k}>0$ be the associated Lipschitz moduli of $f$ and $\nabla h_{1}$ respectively. In the majorization stage, we introduce the surrogate function $\mathcal{M}(\cdot\mid x^{k},c_{k})$ of $\widetilde{F}(\cdot,c_{k})$ over $\mathcal{B}(x^{k},\delta_{k})$, which is in the form of
(35) $\mathcal{M}(x\mid x^{k},c_{k})=(2c_{k}-c_{k}^{2}g(x^{k}))f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\left(c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}\right)\|x-x^{k}\|_{2}^{2}+h(x^{k})+\iota_{\mathcal{C}}(x).$

### v2

@@ 3 The connections between problem (1) and the min-max reformulation (10)
$F(x):=\begin{cases}\frac{f(x)}{g(x)}+h(x),&\text{if $x\in\Omega\cap\mathcal{C}$},\\ +\infty,&\text{else.}\end{cases}$ (14)
Obviously, problem (1) can be rewritten as $\min\{F(x):x\in\mathbb{R}^{n}\}$. Recall that $\widetilde{F}$ denotes the objective of the min-max problem (10), as defined in Section 1. By a direct computation, we obtain that for all $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$,
$F(x)-\widetilde{F}(x,c)=f(x)g(x)\left(\frac{1}{g(x)}-c\right)^{2}.$ (15)
@@
$x^{\star}\in\arg\min\left\{\max_{c\in\mathbb{R}}\widetilde{F}(x,c):x\in\Omega\cap\mathcal{C}\right\}.$ (17)
In view of (16), we derive that $1/g(x^{\star})\in\arg\max\{\widetilde{F}(x^{\star},c):c\in\mathbb{R}\}$, which along with (17) indicates that $(x^{\star},1/g(x^{\star}))$ is a global min-max point of (10), according to Definition 2.
Conversely, if $(x^{\star},c_{\star})$ is a global min-max point of (10), then we derive from Definition 2 that $x^{\star}$ satisfies (17) . Thus, we conclude that $x^{\star}\in\arg\min\{F(x):x\in\Omega\cap\mathcal{C}\}$ from (17) and (16). This completes the proof. $\square$
We next establish the connection between the local minimizers of problem (1) and the local min-max points of problem (10).
Proposition 3.2
@@
Proof
We first prove that the statement that $(x^{\star},1/g(x^{\star}))$ is a local min-max point of problem (10) follows from the $x^{\star}$ being a local minimizer of problem (1). Since (16) indicates that $1/g(x^{\star})$ is a global maximizer of $\widetilde{F}(x^{\star},\cdot)$, it suffices to show that, for an arbitrarily fixed $\epsilon>0$, the $x^{\star}$ is a local minimizer of
$\varphi_{\epsilon}(x):=\max\left\{\widetilde{F}(x,c):c\in\mathbb{R},~|c-1/g(x^{\star})|\leq\epsilon\right\},$ (18)
relative to $\Omega\cap\mathcal{C}$ according to Definition 3. Let this $\epsilon>0$ be fixed. Firstly, due to the continuity of $1/g$ around $x^{\star}\in\Omega$, there exists $\delta_{\epsilon}>0$ such that $|1/g(x)-1/g(x^{\star})|\leq\epsilon$ for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})$. In view of this, we deduce from the definition (18) that
$\widetilde{F}(x,1/g(x))\leq\varphi_{\varepsilon}(x)\text{,~~ for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})\text{. }$ (19)
@@
$F(x^{\star})\leq F(x)\text{,~~ for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon}^{\prime})\cap(\Omega\cap\mathcal{C})\text{. }$ (20)
@@
$\varphi_{\epsilon}(x^{\star})\stackrel{{\scriptstyle{(I)}}}{{\leq}}\max_{c\in\mathbb{R}}\widetilde{F}(x^{\star},c)\stackrel{{\scriptstyle{(II)}}}{{=}}F(x^{\star})\stackrel{{\scriptstyle{(III)}}}{{\leq}}F(x)\stackrel{{\scriptstyle{(IV)}}}{{=}}\widetilde{F}(x,1/g(x))\stackrel{{\scriptstyle{(V)}}}{{\leq}}\varphi_{\varepsilon}(x),$ (21)
@@
Conversely, we let $(x^{\star},c_{\star})$ be a local min-max point of (10). Then, according to Definition 3, we know that $c_{\star}$ is a local maximizer of $\widetilde{F}(x^{\star},\cdot)$. Hence, we have $0=\nabla(\widetilde{F}(x^{\star},\cdot))(c_{\star})$, which leads to $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. In addition, also according to Definition 3, we know that $x^{\star}$ is a local minimizer of $\varphi_{\epsilon}$ with some $\epsilon>0$ relative to $\Omega\cap\mathcal{C}$. Thus, there exists some $\delta_{\epsilon}>0$ such that
$\varphi_{\epsilon}(x^{\star})\leq\varphi_{\epsilon}(x),\text{\quad for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$.}$ (22)
@@
$F(x^{\star})=\widetilde{F}(x^{\star},c_{\star})\leq\varphi_{\epsilon}(x^{\star}),$ (23)
where the first equation follows from (16) and $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. As to the right side of (22), it is clear that
$\varphi_{\epsilon}(x)\leq\max\left\{\widetilde{F}(x,c):c\in\mathbb{R}\right\}=F(x),\quad\text{for all $x\in\Omega\cap\mathcal{C}$.}$ (24)
Together with (23) and (24), we deduce from (22) that $F(x^{\star})\leq F(x)$ for any $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$, which yields that $x^{\star}$ is a local minimizer of (1). $\square$
Now we study the relationship between stationary points of problems (1) and (10). To this end, for $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$, we let $\widehat{\partial}_{x}\widetilde{F}(x,c)$ and $\nabla_{c}\widetilde{F}(x,c)$ denote the partial Fréchet subdifferential of $\widetilde{F}$ with respect to $x$ and the partial derivative of $\widetilde{F}$ with respect to $c$, respectively.
Proposition 3.3
@@
$\left\{\begin{aligned} 0\in\widehat{\partial}_{x}\,\widetilde{F}(x^{\star},c_{\star}),\\ 0=\nabla_{c}\widetilde{F}(x^{\star},c_{\star}).\end{aligned}\right$. (25)
@@
Proof
Let $(x^{\star},c_{\star})\in(\Omega\cap\mathcal{C})\times\mathbb{R}$. Due to the continuity of $g$ around $x^{\star}\in\Omega$, there exists $\delta>0$ such that $g(x)>0$ holds for all $x\in\mathcal{B}(x^{\star},\delta)$. Then, we have
$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})=\widehat{\partial}\left(f/g+h+\iota_{\mathcal{C}}-\varphi\right)(x^{\star})=\widehat{\partial}(F-\varphi)(x^{\star}),$ (26)
@@
$\varphi(x)=f(x)g(x)\left(\frac{1}{g(x)}-c_{\star}\right)^{2},\text{~~~ for all $x\in\mathcal{B}(x^{\star},\delta)$}.$ (27)
@@
Conversely, we suppose that (25) holds. Then $0=\nabla_{c}\widetilde{F}(x^{\star},c_{\star})$ implies that $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. Plugging this into (27) and noting that $f$ and $g$ are non-negative, we deduce that $\varphi(x^{\star})=0$ is the minimum of $\varphi$ on $x\in\mathcal{B}(x^{\star},\delta)$. Hence, by invoking the generalized Fermat’s Rule, we derive $0\in\widehat{\partial}\varphi(x^{\star})$. With the help of (Mordukhovich:2006Frechet_subdifferential, , Theorem 3.1), we have
$\widehat{\partial}(F-\varphi)(x^{\star})\subseteq\bigcap_{v^{\star}\in\widehat{\partial}\varphi(x^{\star})}(\widehat{\partial}F(x^{\star})-v^{\star})\subseteq\widehat{\partial}F(x^{\star})-\{0\}.$ (28)
Together with the first relation of (25), the equation (26), and the relation (28), we finally conclude that $0\in\widehat{\partial}F(x^{\star})$. This completes the proof. $\square$
@@
$0\in\widehat{\partial}(c_{\star}f+\iota_{\mathcal{C}})\left(x^{\star}\right)-c_{\star}^{2}f(x^{\star})\partial g\left(x^{\star}\right)+\nabla h_{1}\left(x^{\star}\right)-\partial h_{2}\left(x^{\star}\right).$ (29)
@@
$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})\overset{{(I)}}{=}\widehat{\partial}(2c_{\star}f-c_{\star}^{2}fg+h+\iota_{\mathcal{C}})(x^{\star})$ (30)
$\overset{{(II)}}{=}\widehat{\partial}(c_{\star}f-c_{\star}^{2}f(x^{\star})g-h_{2}+\iota_{\mathcal{C}})(x^{\star})+\nabla\varphi_{2}(x^{\star})+\nabla h_{1}(x^{\star})$
@@
where (I){(I)} follows from the definition of $\widetilde{F}$ (see (10)) and the open $\Omega$, (II){(II)} holds because $\varphi_{2}$ and $h_{1}$ are differentiable at $x^{\star}$, and (III){(III)} holds thanks to (Mordukhovich:2006Frechet_subdifferential, , Theorem 3.1), $c_{\star}^{2}f(x^{\star})\geq 0$, and the real-valued convexity of the functions $g$ and $h_{2}$. Especially, (III){(III)} in (30) becomes an equality when $g$ and $h_{2}$ are differentiable at $x^{\star}$. Invoking (30) and Proposition 3.3, we then obtain the desired results. $\square$
In general, it is impractical to find an $x^{\star}$ satisfying $0\in\widehat{\partial}F(x^{\star})$. Inspired by Corollary 3.4, we introduce a weaker definition of a critical point of $F$ as follows, which is more amenable to algorithmic analysis and computation.
Definition 4
@@
Using the proximity operator of $f_{\mathcal{C}}:=f+\iota_{\mathcal{C}}$, one can derive that $\bar{x}$ is a critical point of $F$ whenever $\bar{x}\in\mathrm{prox}_{\alpha\bar{c}f_{\mathcal{C}}}(\bar{x}-\alpha(\nabla h_{1}(\bar{x})-\bar{z}-\bar{c}^{2}f(\bar{x})\bar{y}))$, or, equivalently
$\mathrm{dist}\left(\bar{x},~\mathrm{prox}_{\alpha\bar{c}f_{\mathcal{C}}}\left(\bar{x}-\alpha(\nabla h_{1}(\bar{x})-\bar{z}-\bar{c}^{2}f(\bar{x})\bar{y})\right)\right)=0$
for some $\alpha>0$, $\bar{c}=1/g(\bar{x})$, and $(\bar{y},\bar{z})\in\partial g(\bar{x})\times\partial h_{2}(\bar{x})$. Thus, for a given $\bar{x}$, the above distance naturally quantifies the violation of the critical point condition (29). Due to this, below we define an $\epsilon$-critical point of $F$ as a point $\bar{x}$ whose violation is at most $\epsilon$.
Definition 5
Given $\epsilon\geq 0$, we say that the $\bar{x}\in\Omega$ is an $\epsilon$-critical point of $F$, if there exist $\alpha>0$ and $(\bar{y},\bar{z})\in\partial g(\bar{x})\times\partial h_{2}(\bar{x})$ such that
$\mathrm{dist}\left(\bar{x},~\mathrm{prox}_{\alpha\bar{c}f_{\mathcal{C}}}\left(\bar{x}-\alpha(\nabla h_{1}(\bar{x})-\bar{z}-\bar{c}^{2}f(\bar{x})\bar{y})\right)\right)\leq\epsilon$ (31)
holds with $\bar{c}=1/g(\bar{x})$ and $f_{\mathcal{C}}:=f+\iota_{\mathcal{C}}$.
In the next section, inspired by the min–max reformulation, we design an alternating maximization proximal descent algorithm (AMPDA) and, under a mild assumption, establish its subsequential convergence to a critical point of $F$ as well as its iteration complexity for finding an $\epsilon$-critical point.
@@ 4 An alternating maximization proximal descent algorithm
In this section, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max reformulation (10), and establish its subsequential convergence to a critical point as well as its iteration complexity for returning an $\epsilon$-critical point of $F$ under a mild assumption.
@@ 4.1 AMPDA and its interpretations
In this subsection, we present AMPDA and then make some interpretations on this algorithm. To this end, we first introduce an auxiliary $Q:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\to(-\infty,+\infty]$, which is defined by
$Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2cf(x)+c^{2}f(x)(g^{*}(y)-\left\langle{x},{y}\right\rangle)+h_{1}(x)+h_{2}^{*}(z)-\left\langle{x},{z}\right\rangle$ (32)
@@
$F(x)\overset{{(I)}}{=}\widetilde{F}(x,c_{x})\overset{{(II)}}{\leq}Q(x,y,z,c_{x}),$ (33)
where (I){(I)} follows from (16), (II){(II)} follows from the Fenchel-Young Inequality and the non-negativity of $c_{x}^{2}f(x)$, and (II){(II)} becomes an equality when $(y,z)\in\partial g(x)\times\partial h_{2}(x)$. Now we formally present our AMPDA in Algorithm 1. Before proceeding, we provide some interpretations and observations on AMPDA. This algorithm can be interpreted as an alternating iteration method for solving the min-max reformulation (10) as illustrated below.
@@
Step 0. Input $x^{0}\in\mathrm{dom}(F)$, $0<\underline{\alpha}\leq\overline{\alpha}$, $\sigma>0$, $0<\gamma<1$, and set $k\leftarrow 0$.
@@
$\widehat{x}^{k}\in\mathrm{prox}_{\alpha c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right).$ (34)
@@
$Q(\widehat{x}^{k},y^{k},z^{k},1/g(\widehat{x}^{k}))+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k});$ (35)
@@
Second, in Step 2 of the $k$-th iteration, AMPDA solves a proximal subproblem with respect to $f_{\mathcal{C}}$, which also requires to evaluate the gradient of $h_{1}$, the subgradients of $g$ and $h_{2}$ at $x^{k}$. As we will see below, this can be viewed as applying one step of the majorization-minimization scheme to the minimization problem $\min\{\widetilde{F}(x,c_{k}):x\in\mathbb{R}^{n}\}$. Since $f$ and $\nabla h_{1}$ are locally Lipschitz continuous on $\Omega$, there exists $\delta_{k}>0$ such that $f$ and $\nabla h_{1}$ are Lipschitz continuous on the neighborhood $\mathcal{B}(x^{k},\delta_{k})\subseteq\Omega$. Let $L_{f,k}>0$ and $L_{\nabla h_{1},k}>0$ be the associated Lipschitz moduli of $f$ and $\nabla h_{1}$ respectively. In the majorization stage, we introduce the surrogate function $\mathcal{M}(\cdot\mid x^{k},c_{k})$ of $\widetilde{F}(\cdot,c_{k})$ over $\mathcal{B}(x^{k},\delta_{k})$, which is in the form of
$\mathcal{M}(x\mid x^{k},c_{k})=(2c_{k}-c_{k}^{2}g(x^{k}))f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\left(c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}\right)\|x-x^{k}\|_{2}^{2}+h(x^{k})+\iota_{\mathcal{C}}(x).$ (36)